OpenStudy (anonymous):

The numbers 7, 24, and c are a Pythagorean triple. What is the value of c? A. 25 B. 27 C. 24 D. 31

OpenStudy (anonymous):

By Pythagoras Theorem, c^2 = 7^2 + 24^2 = 49 + 576 = 625 c = square root of 625 = 25

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I see how you got that, Thank you :D